rewrite this as a function of x
4x-6y= - 30

Answer by praseenakos@yahoo.com(507)

QUESTION:
rewrite this as a function of x
4x-6y= - 30
ANSWER:
4x-6y= - 30
Subtract 4x from both sides of the expression
==> 4x-6y -4x = - 30 - 4x
==>4x-4x -6y = - 30 - 4x
==> -6y = - 30 - 4x
Now divide whole expression by -6
==> -6y/-6 = - 30/-6 - 4x/-6
==> y = 5 + (4/6)x
which is a function of x.
Regards.
Praseena.
praseenakos@yahoo.co.in